OpenStudy (anonymous):

I need help proving these trig equations sin(x)csc(x)+ 1/(1-sin^2(x))=secx (1+tan^2(x))/(1-tan^2(x)=sin(x)-cos(x)

hero (hero):

So the first one, the setup is: \[\sin(x)\csc(x) + \frac{1}{1 - \sin^2(x)} = \sec(x)\]

OpenStudy (anonymous):

Yes

hero (hero):

Are you 100% sure about the way I setup the first one?

OpenStudy (anonymous):

Yea im 96-98% sure it is that way but what im unsure of is how to get 1/cos or 1+tan^2x

hero (hero):

If it's not true, then there's no point in trying to verify it.

OpenStudy (anonymous):

Alright Thank you for your help :)

hero (hero):

Hang on for a minute. Where are you going?

OpenStudy (anonymous):

nowhere i guess :o?

hero (hero):

Okay, I'm going to show you a couple things

OpenStudy (anonymous):

Alright

hero (hero):

Well, the explanation is more complicated than I thought.

OpenStudy (anonymous):

alright :o

hero (hero):

What was the exact wording of the original question?

OpenStudy (anonymous):

Prove the equations 1 step at a time

hero (hero):

Okay, well, basically, I would begin like this: \[\sin(x) \dot\ \frac{1}{\sin(x)} + \frac{1}{1 - \sin^2(x)} = \sec(x)\]

hero (hero):

sin(x) and csc(x) are inverses of each other so they cancel to just 1 so what you really have is: \[1 + \frac{1}{1 - \sin^2(x)} = \sec(x)\]

hero (hero):

Which can be re-written as: \[\frac{1 - \sin^2x}{1 - \sin^2x} + \frac{1}{1 - \sin^2x} = \sec(x)\] and combined to get: \[\frac{1 - \sin^2x + 1}{1 - \sin^2x} = \sec(x)\] Which simplifies to just: \[\frac{- \sin^2x}{1 - \sin^2x} = \sec(x)\] or \[-\frac{\sin^2x}{1 - \sin^2x} = \sec(x)\]

hero (hero):

Now we can replace 1 with the pythagorean identity: \(\sin^2x + \cos^2x = 1\) \[-\frac{\sin^2x}{\sin^2x + \cos^2x - \sin^2x} = \sec(x)\] Which simplifes to : \[-\frac{\sin^2x}{\cos^2x} = \sec(x)\] or \[-\tan^2x= \sec(x)\]

hero (hero):

Now we know this is not true because there's no way that: \[-\frac{\sin^2x}{\cos^2x} = \frac{1}{\cos(x)}\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ive gone at it with these 2 problems for 3 hours with the result of no solution being something completely out of the question to me

hero (hero):

Even if you multiplied both sides by \(\cos^2x\) to get: \(-\sin^2x = \cos(x)\) At that point, you could graph \(\sin^2x\) and \(\cos(x)\) and visually see that they are not equal

OpenStudy (anonymous):

yep

hero (hero):

You basically do a similar process with the second one and then realize that one isn't true either.

OpenStudy (anonymous):

Thank you :)

hero (hero):

To save yourself from all the trouble, you should get a graphing calc.

OpenStudy (anonymous):

Yea but on tests we cant use graphing calcs

hero (hero):

Sorry to hear that

OpenStudy (anonymous):

Yea You have a good afternoon or night depending on where your from :)

hero (hero):

You too